OpenStudy (anonymous):

PLEASE EXPLAIN STEP BY STEP find the remaining 5 trig functions when sec(theta)=-2 and cot(theta)>0

OpenStudy (freckles):

if cos is negative and cot is positive, then sin is ?

OpenStudy (anonymous):

negative.

OpenStudy (anonymous):

but how do i find the other five functions with the info given?

OpenStudy (freckles):

cool so since cos is neg and sin is neg then we are in the third quadrant

OpenStudy (freckles):

we will build a triangle there

OpenStudy (freckles):



OpenStudy (freckles):

use that sec(theta)=-2 to label the triangle

OpenStudy (freckles):

sec(theta)=hyp/adj =2/(-1)

OpenStudy (freckles):

you will need to find the opp side measurement by use of Pythagorean theorem

OpenStudy (freckles):

then the rest is just using the definition of the trig functions

OpenStudy (freckles):

\[\sin(\theta)=\frac{opp}{hyp} \\ \cos(\theta)=\frac{adj}{hyp} \\ \tan(\theta)=\frac{opp}{adj} \\ \csc(\theta)=\frac{hyp}{opp} \\ \sec(\theta)=\frac{hyp}{adj} \\ \cot(\theta)=\frac{adj}{opp}\]